OpenStudy (anonymous):

Show an equation and a solution for the problem. At noon, a train leaves Bridgetown heading east at 90 mph to Cartertown. Another train leaves Bridgetown at 1:00 p.m., also heading east at 100 mph. In how many hours after the second train left will they pass each other?

OpenStudy (lilg132):

9 hours

OpenStudy (lilg132):

speed of first train = A speed of second train = B Extra amount of miles being covered by second train = E Amount of time taken to pass = T B - A = E T = A / E

OpenStudy (lilg132):

so that will equal to 100 - 90 = 10 90 / 10 = 9 Hours = T